# Filling in the Blanks:

### v1

@@ Abstract
Aims. To present and test a method to infer a probability density function (PDF) for the missing $v_{\mathrm{los}}$ of a star with 5D information within $2.5\,\mathrm{kpc}$. We use stars from the Gaia DR3 RVS catalogue to describe the local orbital structure in action space. This technique also allows us to infer the probability that a 5D star is associated with the Milky Way’s stellar Disc or the stellar Halo, which can be further decomposed into known stellar substructures.
Methods. The method is tested on a 6D Gaia DR3 RVS sample and a 6D Gaia sample crossmatched to ground-based spectroscopic surveys, stripped of their true $v_{\mathrm{los}}$. The stars predicted $v_{\mathrm{los}}$, membership probabilities, and inferred structure properties are then compared to the true 6D equivalents, allowing the method’s accuracy and limitations to be studied in detail.
@@ 1 Introduction
Typically, dynamically detecting individual substructures requires large samples of stars with accurate kinematics (Helmi & de Zeeuw, 2000; Helmi, 2020). The Gaia mission (Gaia Collaboration et al., 2016) has convincingly delivered such data, substantially enriching our understanding of the MW and its assembly history. Following the release of Gaia DR2 (Gaia-Collaboration et al., 2018a), it was confirmed that the inner stellar halo is dominated by a single massive accreted dwarf galaxy, Gaia Enceladus Sausage (GES) (Belokurov et al., 2018; Helmi et al., 2018). Soon after, further halo substructure was found in the Gaia dataset such as Sequoia (Myeong et al., 2019), Thamnos (Koppelman et al., 2019b), among others (Yuan et al., 2018; Myeong et al., 2022). The most recent Gaia data release, DR3 (Gaia Collaboration et al., 2023), has been found to contain even more substructure. Works such as Dodd et al. (2023), hereafter Dodd23, identify new dynamical groups labelled ED-1 to ED-6, alongside other recent discoveries such as Typhon (Tenachi et al., 2022).
@@
These limitations associated with working within the RVS catalogue could be avoided if the lack of $v_{\mathrm{los}}$ could be overcome or mitigated. Some works have identified likely halo stars in the 5D dataset using cuts in transverse sky velocity (Gaia Collaboration et al., 2018b), or defined samples based on reduced proper motions (Koppelman & Helmi, 2021; Viswanathan et al., 2023). These 5D halo datasets can then be used by techniques such as the STREAMFINDER algorithm, which can find coherent stellar streams upon the sky using 5D information (Ibata et al., 2021). For specific phase-mixed structures, the effect of missing $v_{\mathrm{los}}$ information can be minimised by focusing on specific regions of the sky, such as the Galactic anticentre. In Koppelman et al. (2019a) and Ruiz-Lara et al. (2022b), this approach is used to identify candidate members of the Helmi Streams (Helmi et al., 1999), which can be distinguished through $\mu_{b}$ alone because of their prominent vertical velocity.
An alternative approach is to attempt to constrain and estimate the missing $v_{\mathrm{los}}$. Using the 6D Gaia RVS sample, statistical relationships between the 5D coordinates and the $v_{\mathrm{los}}$ can be built, allowing the likely $v_{\mathrm{los}}$ of a star to be inferred. In Mikkola et al. (2023) (hereafter Mikkola23), a penalised maximum-likelihood method (Dehnen, 1998; Mikkola et al., 2022) is used to infer full velocity distributions of the 5D Gaia sample. This methodology reveals features in velocity space hidden in the 5D sample, which can then be used to estimate the number of stars expected to belong to specific stellar substructures. Similarly, Naik & Widmark (2024) (hereafter Naik24) is the latest in a series of papers (Dropulic et al., 2021, 2023) that use Bayesian neural networks (Titterington, 2004) trained on a 6D Gaia RVS sample to predict the missing $v_{\mathrm{los}}$ of a 5D Gaia sample. The predictions are probability distributions with a typical uncertainty of $25$-$30$ $\,\text{km}\hskip 1.00006pt\text{s}^{-1}$.
@@ 2 Data
Figure 1: A local sample of Gaia DR3 stars in different dynamical spaces (panels), classified into different groups (indicated by colour) by Dodd23. The panels show the spaces of energy and the vertical and perpendicular components of angular momentum used in Dodd23 to identify the Halo groups. The Disc group (purple) is kinematically selected with a Toomre velocity cut; the rest of the structures are found in the complementary kinematically selected Halo. In the plot, scatter points representing the Disc group have been down-sampled by a factor of 500, whilst the population in the legend is the total population of Disc stars.
@@
To convert to Galactocentric coordinates, we assume a solar motion of ${\left(U,V,W\right)}_{\odot}=(11.1,12.24,7.25)\,\,\text{km}\hskip 1.00006pt\text{s}^{-1}$ (Schönrich et al., 2010), a local standard of rest of $\left|\textbf{V}_{\textrm{LSR}}\right|=232.8\,\text{km}\hskip 1.00006pt\text{s}^{-1}$ (McMillan, 2017), and Solar radius of $R_{\odot}=8.2\,\mathrm{kpc}$ (McMillan, 2017). We use the potential of Cautun et al. (2020), which differs from Dodd23, but we find no substantial effect on the groupings of Dodd23. For dynamical calculations, we use the software package agama (Vasiliev, 2019), which uses the Stäckel fudge algorithm to calculate actions (Binney, 2012; Sanders & Binney, 2016).
@@ 2.1 Stellar Halo Groups
The Halo sample is decomposed into dynamical groups identified in Dodd23. This work built on the previous studies of Lövdal et al. (2022) and Ruiz-Lara et al. (2022a), applying a clustering algorithm to identify statistically significant overdensities in the space of energy and angular momentum components to find clusters of stars in their Halo sample. These clusters were then joined to form groups using a single linkage algorithm and a distance cut based on the dendrogram derived by the algorithm. Finally, stars located within a Mahalanobis distance of $2.2$ from each group’s centre (which approximately corresponds to a contour containing 80% of the original members) have been added to the groups to include additional members. These groups are thus likely to be conservative, although not necessarily pure (such as Thamnos; see Dodd et al., 2022).
We use the original groups of Dodd23 with a few minor modifications. Several “loose” clusters [^2] Clusters 34, 48, and 77 in Dodd23 that fall within the Mahalanobis distance of GES but fall slightly over the dendrogram cut used in Dodd23 to define the group are joined to GES. We choose not to include the LRL3 group, as it is a low-energy structure of mixed populations with an unclear origin. Instead, its’ stars are classified as Ungrouped Halo. We believe that LRL64 is the same as the structure Antaeus identified in Oria et al. (2022), and we chose to use the original name designated in Ruiz-Lara et al. (2022a). The membership of ED-2 is updated to members identified in Balbinot et al. (2023), increasing from 32 to 34 members within our RVS sample. Similarly, the membership of Typhon is updated to members identified in Tenachi et al. (2022). The final groups used are GES, Hot Thick Disc (HTD), Thamnos, Helmi Streams (HS), Sequoia, LRL64, ED-2, ED-3, ED-4, ED-5, ED-6, and Typhon. Their exact populations can be found in the legend of Fig. 1. This classification leaves approximately 57,000 Ungrouped Halo stars, which are treated as another group.
@@ 3.1 Selection Function Correction
$F_{\odot}\left(\mathbf{J}\right)=F\left(\mathbf{J}\right)\,\,\forall\,\,\bf{J}\,\,\text{that cross the Solar neighbourhood}$ (3)
@@
$F_{\mathrm{obs}}\left(\mathbf{J}\right)\approx\tau\left(\bm{J}\right)\ F_{\odot}\left(\mathbf{J}\right)$ (4)
@@
First, for an orbit $\bf{J}$, we sample many physical points along its trajectory. Next, we define a typical halo stellar population with an isochrone, using a Chabrier (Chabrier, 2003) Initial Mass Function and an alpha-enhanced BaSTI isochrone (Pietrinferni et al., 2021), with a metallicity of $\left[\mathrm{Fe}/\mathrm{H}\right]=-1.1$ and an age of $11$ Gyrs. Then, the probability of observing a star of a given magnitude and colour from the stellar population at the sampled positions is evaluated. This probability is estimated using the dust maps of Lallement et al. (2022), photometric corrections of Gaia Collaboration et al. (2018b), and the RVS selection function of Castro-Ginard et al. (2023). Observational uncertainties and the corresponding quality cuts are estimated using the PyGaia [^4] https://github.com/agabrown/PyGaia software package. Finally, by averaging over the defined halo stellar population and over the points around the orbit, the expected observable fraction $\tau\left(\bm{J}\right)$ of stars upon the orbit $\bf{J}$ is found.
@@ 4 Method
Figure 2: Our sample of 6D stars in action space, classified into different groups of Dodd23 (indicated by colour). The Ungrouped Halo stars are represented by smoothed density, as indicated by the colours and contours in the colour bar. This density, and those of the individual groups, are corrected for selection effects (see text). Across the space is an example path (red line) of a 5D star through action space found by varying an assumed $v_{\mathrm{los}}$, with values indicated in the legend. The corresponding PDF can be seen in the following Fig. 3. This example star is a GES member from the 6D Gaia RVS sample, with the true $v_{\mathrm{los}}$ and the corresponding point in action space marked by a square.
Figure 3: A PDF of $v_{\mathrm{los}}$, calculated from a 5D star’s path in action space (see Fig. $2$). The thick black line total PDF shows the total density. This total can be decomposed into contributions from different groups, denoted by different colours. The probability that the star belongs to these different groups can be found by integrating the area under the respective lines, shown in the legend. As a star from the 6D sample, the true $v_{\mathrm{los}}$ is marked in green, and it is known that it is a GES member. Our method finds a $63\%$ probability of belonging to GES but could also belong to Ungrouped Halo or Sequoia.
@@
$p\left(v_{\mathrm{los}}|\bm{u}\right)\propto F\left(\bm{J},\bm{\theta}\right)$ (9)
@@
$p\left(v_{\mathrm{los}}|\bm{u}\right)=\frac{1}{\lambda\left(\bm{u}\right)}F\left(\bm{J},\bm{\theta}\right)|_{\left(v_{\mathrm{los}},\bm{u}\right)}$ (10)
@@
$\lambda\left(\bm{u}\right)=\int F\left(\bm{J},\bm{\theta}\right)|_{\left(v_{\mathrm{los}},\bm{u}\right)}\operatorname{d}v_{\mathrm{los}}$ (11)
@@
$p\left(v_{\mathrm{los}}|\bm{u}\right)\propto F_{\mathrm{obs}}^{*}\left(\mathbf{J}\right)$ (12)
@@ 4.1 Group membership
$p_{g}\left(v_{\mathrm{los}}|\bm{u}\right)=\frac{1}{\lambda\left(\bm{u}\right)}F_{g}\left(\bm{J},\bm{\theta}\right)|_{v_{\mathrm{los}},\bm{u}}$ (14)
@@ 4.2 Phase-Clumped and Partially Phase-Mixed Substructures
Figure 4: The stars of un-phase-mixed groups in radial and vertical velocity space (in Galactic cylindrical coordinates). If the groups were phase-mixed, there would be a roughly equal number of stars from each group in every quadrant. Whilst not every group is in a single quadrant, none are consistent with being evenly distributed.
@@
The velocity length scale $\sigma_{v}$ is estimated for each group by considering the largest range $\frac{1}{2}\left(84\%-16\%\right)$ of the distribution of different absolute velocity components, and can be seen in the caption of Fig. 4. These length scales are deliberately chosen to be generous so as not to be overly restrictive while effectively describing the space. These velocity criteria are applied as an additional selection function when calculating the correction weight for stars in the phase-clumped groups, effectively reducing the volume of angle-space that the star can be observed in and increasing the correction weight.
@@
$F_{g}\left(\bm{J},\bm{\theta}\right)|_{v_{\mathrm{los}},\bm{u}}\approx\frac{1}{{\left(2\pi\right)}^{3}}\eta_{g}\left(v_{R},v_{Z},v_{\phi}\right)F^{*}_{g}\left(\bm{J}\right)|_{v_{\mathrm{los}},\bm{u}}$ (17)
@@
5 Testing $v_{\mathrm{los}}$ Predictions
Figure 5: The distribution of CDF values $H\left(v_{\mathrm{True}}\right)$ of the true $v_{\mathrm{los}}$ of the stars in the predicted PDFs of each star in our Gaia RVS sample. If all assumptions and the PDFs are correct, these values should be distributed uniformly. We plot the total distribution of all stars and the distributions of the Disc and Halo stars (with CDF values corresponding to their respective PDFs). The total statistical error $\Delta_{\%}$ of our PDFs can be estimated by integrating the difference between the found and true uniform distribution.

### v2

@@ Abstract
Aims. We aim to present and test a method to infer a probability density function (PDF) for the missing $v_{\mathrm{los}}$ of a star with 5D information within $2.5\,\mathrm{kpc}$. This technique also allows us to infer the probability that a 5D star is associated with the Milky Way’s stellar Disc or the stellar Halo, which can be further decomposed into known stellar substructures.
Methods. We use stars from the Gaia DR3 RVS catalogue to describe the local orbital structure in action space. The method is tested on a 6D Gaia DR3 RVS sample and a 6D Gaia sample crossmatched to ground-based spectroscopic surveys, stripped of their true $v_{\mathrm{los}}$. The stars predicted $v_{\mathrm{los}}$, membership probabilities, and inferred structure properties are then compared to the true 6D equivalents, allowing the method’s accuracy and limitations to be studied in detail.
@@ 1 Introduction
Typically, dynamically detecting individual substructures requires large samples of stars with accurate kinematics (Helmi & de Zeeuw, 2000; Helmi, 2020). The Gaia mission (Gaia Collaboration et al., 2016) has convincingly delivered such data, substantially enriching our understanding of the MW and its assembly history. Following the release of Gaia DR2 (Gaia-Collaboration et al., 2018), it was confirmed that the inner stellar halo is dominated by a single massive accreted dwarf galaxy, Gaia Enceladus Sausage (GES) (Belokurov et al., 2018; Helmi et al., 2018). Soon after, further halo substructure was found in the Gaia dataset such as Sequoia (Myeong et al., 2019), Thamnos (Koppelman et al., 2019a), among others (Yuan et al., 2018; Myeong et al., 2022). The most recent Gaia data release, DR3 (Gaia Collaboration et al., 2023), has been found to contain even more substructure. Works such as Dodd et al. (2023), hereafter Dodd23, identify new dynamical groups labelled ED-1 to ED-6, alongside other recent discoveries such as Typhon (Tenachi et al., 2022).
@@
These limitations associated with working within the RVS catalogue could be avoided if the lack of $v_{\mathrm{los}}$ could be overcome or mitigated. Some works have identified likely halo stars in the 5D dataset using cuts in transverse sky velocity (Gaia Collaboration et al., 2018), or defined samples based on reduced proper motions (Koppelman & Helmi, 2021; Viswanathan et al., 2023). These 5D halo datasets can then be used by techniques such as the STREAMFINDER algorithm, which can find coherent stellar streams upon the sky using 5D information (Ibata et al., 2021). For specific phase-mixed structures, the effect of missing $v_{\mathrm{los}}$ information can be minimised by focusing on specific regions of the sky, such as the Galactic anticentre. In Koppelman et al. (2019b) and Ruiz-Lara et al. (2022a), this approach is used to identify candidate members of the Helmi Streams (Helmi et al., 1999), which can be distinguished through $\mu_{b}$ alone because of their prominent vertical velocity.
An alternative approach is to attempt to constrain and estimate the missing $v_{\mathrm{los}}$. Using the 6D Gaia RVS sample, statistical relationships between the 5D coordinates and the $v_{\mathrm{los}}$ can be built, allowing the likely $v_{\mathrm{los}}$ of a star to be inferred. In Mikkola et al. (2023) (hereafter Mikkola23), a penalised maximum-likelihood method (Dehnen, 1998; Mikkola et al., 2022) is used to infer full velocity distributions of the 5D Gaia sample. This methodology reveals features in velocity space hidden in the 5D sample, which can then be used to estimate the number of stars expected to belong to specific stellar substructures. Similarly, Naik & Widmark (2024) (hereafter Naik24) is the latest in a series of papers (Dropulic et al., 2021, 2023) that use neural networks trained on a 6D Gaia RVS sample to predict the missing $v_{\mathrm{los}}$ of a 5D Gaia sample. The work of Naik24, uses Bayesian neural networks (Titterington, 2004) to infer predicted probability distributions with a typical uncertainty of $25$-$30$ $\,\text{km}\hskip 1.00006pt\text{s}^{-1}$.
@@ 2 Data
Figure 1: Local sample of Gaia DR3 stars in different dynamical spaces (panels), classified into different groups (indicated by colour) by Dodd23. The panels show the spaces of energy and the vertical and perpendicular components of angular momentum used in Dodd23 to identify the Halo groups. The Disc group (purple) is kinematically selected with a Toomre velocity cut; the rest of the structures are found in the complementary kinematically selected Halo. In the plot, scatter points representing the Disc group have been down-sampled by a factor of 500, whilst the population in the legend is the total population of Disc stars.
@@
To convert to Galactocentric coordinates, we assume a solar motion of ${\left(U,V,W\right)}_{\odot}=(11.1,12.24,7.25)\,\,\text{km}\hskip 1.00006pt\text{s}^{-1}$ (Schönrich et al., 2010), a local standard of rest of $\left|\textbf{V}_{\textrm{LSR}}\right|=232.8\,\text{km}\hskip 1.00006pt\text{s}^{-1}$ (McMillan, 2017), and Solar radius of ${{\color[rgb]{0,0,0}R_{0}}}=8.2\,\mathrm{kpc}$ (McMillan, 2017). We use the potential of Cautun et al. (2020), which differs from Dodd23, but we find no substantial effect on the groupings of Dodd23. For dynamical calculations, we use the software package agama (Vasiliev, 2019), which uses the Stäckel fudge algorithm to calculate actions (Binney, 2012; Sanders & Binney, 2016).
@@ 2.1 Stellar halo groups
The Halo sample is decomposed into dynamical groups identified in Dodd23. This work built on the previous studies of Lövdal et al. (2022) and Ruiz-Lara et al. (2022b), applying a clustering algorithm to identify statistically significant overdensities in the space of energy and angular momentum components to find clusters of stars in their Halo sample. These clusters were then joined to form groups using a single linkage algorithm and a distance cut based on the dendrogram derived by the algorithm. Finally, stars located within a Mahalanobis distance of $2.2$ from each group’s centre (which approximately corresponds to a contour containing 80% of the original members) have been added to the groups to include additional members. These groups are thus likely to be conservative, although not necessarily pure (such as Thamnos; see Dodd et al., 2022).
We use the original groups of Dodd23 with a few minor modifications. Several “loose” clusters [^2] Clusters 34, 48, and 77 in Dodd23 that fall within the Mahalanobis distance of GES but fall slightly over the dendrogram cut used in Dodd23 to define the group are joined to GES. We choose not to include the LRL3 group, as it is a low-energy structure of mixed populations with an unclear origin. Instead, its’ stars are classified as Ungrouped Halo. We believe that LRL64 is the same as the structure Antaeus identified in Oria et al. (2022), and we chose to use the original name designated in Ruiz-Lara et al. (2022b). The membership of ED-2 is updated to members identified in Balbinot et al. (2023), increasing from 32 to 34 members within our RVS sample. Similarly, the membership of Typhon is updated to members identified in Tenachi et al. (2022). The final groups used are GES, Hot Thick Disc (HTD), Thamnos, Helmi Streams (HS), Sequoia, LRL64, ED-2, ED-3, ED-4, ED-5, ED-6, and Typhon. Their exact populations can be found in the legend of Fig. 1. This classification leaves approximately 57,000 Ungrouped Halo stars, which are treated as another group.
@@ 3.1 Selection function correction
$F_{\odot}\left(\mathbf{J}\right)=F\left(\mathbf{J}\right)\,\,\forall\,\,\bf{J}\,\,\text{that cross the Solar neighbourhood}.$ (3)
@@
$F_{\mathrm{obs}}\left(\mathbf{J}\right)\approx\tau\left(\bm{J}\right)\ F_{\odot}\left(\mathbf{J}\right).$ (4)
@@
First, for an orbit $\bf{J}$, we sample many physical points along its trajectory. Next, we define a typical halo stellar population with an isochrone, using a Chabrier (Chabrier, 2003) Initial Mass Function and an alpha-enhanced BaSTI isochrone (Pietrinferni et al., 2021), with a metallicity of $\left[\mathrm{Fe}/\mathrm{H}\right]=-1.1$ and an age of $11$ Gyrs. Then, the probability of observing a star of a given magnitude and colour from the stellar population at the sampled positions is evaluated. This probability is estimated using the dust maps of Lallement et al. (2022), photometric corrections of Gaia Collaboration et al. (2018), and the RVS selection function of Castro-Ginard et al. (2023). Observational uncertainties and the corresponding quality cuts are estimated using the PyGaia [^4] https://github.com/agabrown/PyGaia software package. Finally, by averaging over the defined halo stellar population and over the points around the orbit, the expected observable fraction $\tau\left(\bm{J}\right)$ of stars upon the orbit $\bf{J}$ is found.
@@ 4 Method
Figure 2: Our sample of 6D Gaia DR3 RVS stars in action space, classified into different groups of Dodd23 (indicated by colour). The Ungrouped Halo stars are represented by smoothed density, as indicated by the colours and contours in the colour bar. This density, and those of the individual groups, are corrected for selection effects (see text). Across the space is an example path (red line) of a 5D star through action space found by varying an assumed $v_{\mathrm{los}}$, with values indicated in the legend. The corresponding PDF can be seen in the following Fig. 3. This example star is a GES member from the 6D Gaia RVS sample, with the true $v_{\mathrm{los}}$ and the corresponding point in action space marked by a square.
Figure 3: PDFs of $v_{\mathrm{los}}$, calculated from a 5D star’s path in action space (see Fig. $2$). The thick black line total PDF shows the total density. This total can be decomposed into contributions from different groups, denoted by different colours. The probability that the star belongs to these different groups can be found by integrating the area under the respective lines, shown in the legend. As a star from the 6D sample, the true $v_{\mathrm{los}}$ is marked in green, and it is known that it is a GES member. Our method finds a $63\%$ probability of belonging to GES but could also belong to Ungrouped Halo or Sequoia.
@@
$p\left(v_{\mathrm{los}}|\bm{u}\right)\propto F\left(\bm{J},\bm{\theta}\right).$ (9)
@@
$p\left(v_{\mathrm{los}}|\bm{u}\right)=\frac{1}{\lambda\left(\bm{u}\right)}F\left(\bm{J},\bm{\theta}\right)|_{\left(v_{\mathrm{los}},\bm{u}\right)},$ (10)
@@
$\lambda\left(\bm{u}\right)=\int F\left(\bm{J},\bm{\theta}\right)|_{\left(v_{\mathrm{los}},\bm{u}\right)}\operatorname{d}v_{\mathrm{los}},$ (11)
@@
$p\left(v_{\mathrm{los}}|\bm{u}\right)\propto F_{\mathrm{obs}}^{*}\left(\mathbf{J}\right).$ (12)
@@ 4.1 Group membership
$p_{g}\left(v_{\mathrm{los}}|\bm{u}\right)=\frac{1}{\lambda\left(\bm{u}\right)}F_{g}\left(\bm{J},\bm{\theta}\right)|_{v_{\mathrm{los}},\bm{u}},$ (14)
@@ 4.2 Phase-clumped and partially phase-mixed substructures
Figure 4: Stars of un-phase-mixed groups in radial and vertical velocity space (in Galactic cylindrical coordinates). If the groups were phase-mixed, there would be a roughly equal number of stars from each group in every quadrant. Whilst not every group is in a single quadrant, none are consistent with being evenly distributed.
@@
The velocity length scale $\sigma_{v}$ is estimated for each group by considering the largest range $\frac{1}{2}\left(84\%-16\%\right)$ of the distribution of different absolute velocity components, and can be seen in the legend of Fig. 4. These length scales are deliberately chosen to be generous so as not to be overly restrictive while effectively describing the space. These velocity criteria are applied as an additional selection function when calculating the correction weight for stars in the phase-clumped groups, effectively reducing the volume of angle-space that the star can be observed in and increasing the correction weight.
@@
$F_{g}\left(\bm{J},\bm{\theta}\right)|_{v_{\mathrm{los}},\bm{u}}\approx\frac{1}{{\left(2\pi\right)}^{3}}\eta_{g}\left(v_{R},v_{Z},v_{\phi}\right)F^{*}_{g}\left(\bm{J}\right)|_{v_{\mathrm{los}},\bm{u}}.$ (17)
@@
5 Testing vlos predictions
Figure 5: Distributions of CDF values $H\left(v_{\mathrm{True}}\right)$, calculated from the true $v_{\mathrm{los}}$ of the stars and the PDFs of our method, for each star in our Gaia RVS sample. If all assumptions and the PDFs are correct, these values should be distributed uniformly. We plot the total distribution of all stars and the distributions of the Disc and Halo stars (with CDF values corresponding to their respective PDFs). The total statistical error $\Delta_{\%}$ of our PDFs can be estimated by integrating the difference between the found and true uniform distribution.